# Model-Agnostic, Temperature-Informed Sampling Enhances Cross-Year Crop Mapping with Deep Learning

### v2

@@ Abstract
Conventional benchmarks for crop type classification from optical satellite time series typically assume access to labeled data from the same year and rely on fixed calendar-day sampling. This limits generalization across seasons, where crop phenology shifts due to interannual climate variability, and precludes real-time application when current-year labels are unavailable. Furthermore, uncertainty quantification is often neglected, making such approaches unreliable for crop monitoring applications. Inspired by ecophysiological principles of plant growth, we propose a simple, model-agnostic sampling strategy that leverages growing degree days (GDD), based on daily average temperature, to replace calendar time with thermal time. By uniformly subsampling time series in this biologically meaningful domain, the method emphasizes phenologically active growth stages while reducing temporal redundancy and noise. We evaluate the method on a multi‐year Sentinel-2 dataset spanning all of Switzerland, training on one growing season and testing on other seasons. Compared to state‐of‐the‐art baselines, our method delivers substantial gains in classification accuracy and, critically, produces more calibrated uncertainty estimates. Notably, our method excels in low-data regimes and enables significantly more accurate early-season classification. With only 10 percent of the training data, our method surpasses the state-of-the-art baseline in both predictive accuracy and uncertainty estimation, and by the end of June, it achieves performance similar to a baseline trained on the full season. These results demonstrate that leveraging temperature data not only improves predictive performance across seasons but also enhances the robustness and trustworthiness of crop‐type mapping in real‐world applications.
keywords:
Crop Mapping, Deep Learning, Transformer, Temperature Data, Time Series, Uncertainty Quantification
@@ 1 Introduction
Monitoring agricultural land use is crucial for the sustainability of our food systems (Gómez et al.,, 2016). As the world’s population grows, climates shift, and diets evolve, there is a pressing need for monitoring tools that deliver reliable, affordable, and scalable insights into farming practices. In contrast to traditional methods, such as field surveys and self-reported data, which can be labor-intensive, expensive, and susceptible to errors, remote sensing-based approaches offer a faster, more cost-effective, and less subjective alternative for tracking land-use dynamics (Weiss et al.,, 2020).
Recent progress in remote sensing technologies and machine learning algorithms has significantly enhanced the ability to monitor agricultural land use. High-resolution, multi-spectral data from platforms such as Sentinel-2 provide frequent, detailed observations of croplands at minimal cost. By applying advanced machine learning methods, including conventional ensemble learners, e.g, XGBoost (Chen and Guestrin,, 2016) and deep neural networks, e.g, Transformers (Vaswani et al.,, 2017), to these time series images, we can capture crop-specific growth dynamics and distinguish among diverse management practices with high precision. These integrated approaches enable automated, scalable mapping of crop types and growth stages, supporting timely, data driven decisions for sustainable agriculture (Turkoglu,, 2023).
Figure 1: Cumulative Growing Degree Days (GDD) across Switzerland for multiple years.
Modern deep learning methods (Garnot and Landrieu, (2021), Turkoglu et al., 2021a , Metzger et al., (2021), Tarasiou et al., (2023), Rußwurm and Körner, (2018)) excel in classifying crop types from optical satellite image time series. However, these models are typically trained on curated datasets with limited size and/or validated on data collected within a single growing season (Rußwurm et al., (2019), Garnot and Landrieu, (2021), Schneider et al., (2023), Turkoglu et al., 2021a ), neglecting interannual shifts in growth driven by climate variability and thus limiting their applicability across years and regions. Furthermore, their dependence on current-year labeled data for training precludes real time deployment during the ongoing season. As also noted by Roscher et al., (2024), model development should focus on datasets with real-world relevance to further improve accuracy, generalization ability, and real impact on end user applications. Lastly, although these approaches may achieve state-of-the-art accuracy, they rarely incorporate uncertainty quantification, leaving predictive confidence unassessed and often uncalibrated. Ensuring the models are well-calibrated makes them reliable components of agricultural monitoring systems, where miscalibration could otherwise lead to costly resource misallocations or risk mismanagement.
Figure 2: NDVI time series of two parcels with the same crop rotation over two years (2021 and 2022), aligned by calendar date (left) and by cumulative growing degree days (GDD) (right).
@@
In this work, we build upon the state-of-the-art attention-based deep learning architecture proposed by Garnot and Landrieu, (2021), and introduce a simple, model-agnostic sampling strategy that incorporates daily average temperature to reduce redundancy and noise in satellite image time series. Deep learning models for crop classification typically rely on subsampled time series, as crop development progresses gradually and many satellite observations are temporally redundant. Subsampling is also necessary to reduce sequence length for efficient model training and inference within limited GPU memory. To better guide this selection, we replace calendar days with growing degree days (GDD) as the temporal reference and perform uniform subsampling in this thermal time domain. This approach emphasizes the crop’s biologically active periods, ensuring that the model focuses on the most informative observations.
@@ 2 Related Work
Deep Learning for Crop Mapping from Satellite Imagery.
Deep learning has dramatically improved crop‐type classification from satellite time series, outperforming traditional methods based on handcrafted features (Turkoglu,, 2023). Early sequence‐modeling approaches used recurrent architectures to capture phenological dynamics. Long short‐term memory (LSTM) networks demonstrated strong performance on multi‐temporal inputs (Rußwurm and Körner,, 2017), and the convolutional‐RNN (convRNN) version further boosted both accuracy and efficiency (Turkoglu et al., 2021a, ).
Attention mechanisms have emerged as a powerful alternative for modeling long‐range temporal dependencies. Rußwurm and Körner, (2020) and Garnot et al., (2020) were the first to apply self‐attention to crop classification from Sentinel-2 data and showed improved results. Series, (2020) then proposed the lightweight Temporal Attention Encoder (L-TAE), and Garnot and Landrieu, (2021) embedded L-TAE within a U-Net backbone to adaptively weigh temporal features across space and time. More recently, fully attention-based Vision Transformer (ViT) adaptations have been explored for crop mapping (Bai et al.,, 2023), and Qin et al., (2024) leveraged the recently proposed Mamba state‐space architecture to improve spatio‐temporal classification performance.
To incorporate domain knowledge in deep learning architectures, several works introduce structured priors and hierarchical labels. Turkoglu et al., 2024b exploited multi‐scale crop class hierarchies within a parameter‐efficient convRNN (i.e, convSTAR (Turkoglu et al., 2021b, )) framework, and prototype‐based supervision has been used to integrate class relationships into metric learning (Sainte Fare Garnot and Landrieu,, 2020). Positional encodings tailored to agricultural cycles have also been proposed: (Nyborg et al.,, 2022) introduced thermal-time encoding, using cumulative growing degree days, to normalize phenological shifts across regions. This is the most relevant research for this paper. Moreover, practical deployment demands handling early-season data and irregular revisit intervals. Rußwurm et al., (2023) added an early classification reward loss to LSTM training, improving accuracy in the season when only partial time series are available. Metzger et al., (2021) leveraged neural ordinary differential equations (NODE) to process noisy and irregularly sampled time series, and Cai et al., (2023)’s collect–update–distribute design further addresses sampling irregularity.
Multi-Year Crop Mapping from Satellite Imagery.
Multi-year crop mapping faces challenges due to temporal domain shifts such as interannual variability in weather, phenological differences, and sensor calibration issues. These factors often degrade the performance of models trained on single-season data when applied across multiple years. Turkoglu et al., 2024a extended the experiments of Turkoglu et al., 2021a to a multi‐year setting, training and testing on data from two different years, and reporting preliminary results. Blickensdörfer et al., (2022) employed a Random Forest classifier using dense Sentinel-2 and Landsat-8 time series, monthly Sentinel-1 composites, and environmental variables to evaluate the relative contributions of optical, radar, and environmental data across three meteorologically distinct years (2017–2019). Dubrovin et al., (2024) demonstrated that reconstructing multi-year Sentinel-2 NDVI time series with Fourier-based fitting and applying Gradient Boosting enables improved crop mapping across different years. Gao et al., (2023) investigated ways to optimize the use of multi-year samples in a within-season crop classification model. Qi et al., (2024) presented a Google Earth Engine workflow that fuses field-sample data with multi-year Sentinel-2 time series to map cropping systems on smallholder farms, achieving consistently robust results across diverse landscapes. Li et al., (2023) benchmarked deep learning segmentation models (e.g., U-Net and SegFormer (Xie et al.,, 2021)) for crop mapping across multiple years, while Sykas et al., (2022) evaluated convLSTM and convSTAR (Turkoglu et al., 2021b, ) networks in a similar multi-year context. Very recently, Khallaghi et al., (2025) shows improved results for cross-year crop mapping combining a U-Net with photometric augmentation, Tversky-Focal loss, and Monte Carlo (MC) Dropout.
Uncertainty Modeling in Deep Learning.
Predictive uncertainty is commonly split into aleatoric uncertainty, which stems from inherent data variability such as class overlap or sensor noise, and epistemic uncertainty, which reflects model weight uncertainty due to insufficient coverage of the input space (Kendall and Gal,, 2017). Exact Bayesian inference in deep networks is generally intractable, leading to approximate methods like variational inference (e.g., Bayes by Backprop,  Blundell et al., 2015), Markov Chain Monte Carlo techniques such as Stochastic Gradient Langevin Dynamics (Welling and Teh,, 2011). However, these approaches often struggle to scale to large architectures (Gustafsson et al.,, 2019). Probabilistic ensembles, particularly deep ensembles, currently represent the gold standard for epistemic uncertainty estimation in deep networks, providing well‐calibrated and robust predictions by aggregating outputs from multiple independently trained models (Lakshminarayanan et al.,, 2017). However, the need to train and store several large networks incurs substantial computational and memory overhead. To address this, implicit ensemble techniques, such as Monte Carlo (MC) Dropout (Gal and Ghahramani,, 2016), Snapshot Ensembles (Huang et al.,, 2017), Masksembles (Durasov et al.,, 2021), and FiLM-Ensemble (Turkoglu et al.,, 2022), approximate ensemble diversity via stochastic weight perturbations or dynamic weight combinations, significantly reducing resource requirements. Despite these gains in efficiency, implicit methods generally underperform full ensembles in both predictive accuracy and uncertainty calibration. More recently, LoRA-Ensemble has been introduced as a parameter‐efficient alternative for transformer‐based architectures, achieving uncertainty estimates competitive with deep ensembles at a fraction of the cost (Mühlematter et al.,, 2024).
Uncertainty Modeling in Satellite Imagery and Crop Mapping
Uncertainty quantification has become a standard component of satellite-based land-cover and crop-mapping workflows, as it provides essential information for risk-aware decision-making in agricultural management and ecosystem monitoring. Generating per-pixel or per-field uncertainty maps alongside deterministic predictions enables practitioners to identify low-confidence areas and tailor downstream analyses accordingly. For example, Lang et al., (2022) employs a deep ensemble of convolutional neural networks to estimate global canopy height from sparse LiDAR samples and multi-sensor satellite imagery, producing uncertainty intervals that reflect both model and data variability. Becker et al., (2023) extends this ensemble approach to country-scale forest structure estimation at 10m resolution, demonstrating that dense uncertainty fields help flag regions where auxiliary field data are most needed.
In the context of crop yield and type mapping, several recent studies have adopted stochastic inference techniques. Tian et al., (2024) introduces an attention-guided multi-level crop network for winter wheat yield estimation at the county scale, applying MC Dropout during inference to derive spatially explicit uncertainty maps of yield forecasts. Similarly, Khallaghi et al., (2025) leverages MC Dropout in a U-Net segmentation framework to quantify per-pixel uncertainty in annual cropland masks across multiple years in Ghana, using the resulting confidence estimates to discard unreliable predictions. Liu et al., (2024) integrates MC Dropout in a ViT–ResNet feature encoder to produce field-level uncertainty scores, filtering out low-confidence labels to improve overall mapping precision. Beyond dropout-based methods, Baumert et al., (2024) proposes a probabilistic crop-type mapping method in which an ensemble of classifiers, each trained with perturbed inputs and hyperparameters on Sentinel-1, Sentinel-2, and Landsat-8 data, generates multiple crop maps over Germany; per-pixel frequency counts across the ensemble yield robust probability distributions and uncertainty metrics. Maleki et al., (2024) refines the 30m cropland data by harnessing its built-in Random Forest confidence layer: low-certainty regions are iteratively reclassified using targeted ancillary datasets, resulting in a more accurate and self-aware mapping product. Collectively, these approaches demonstrate the critical role of uncertainty modeling in improving the reliability and interpretability of satellite-driven crop and canopy mapping systems.
@@ 3 Method
Our method builds on the U-Net with Temporal Attention Encoder (U-TAE) (Ronneberger et al.,, 2015; Garnot and Landrieu,, 2021), a state-of-the-art deep learning model for dense segmentation of satellite image time series. Prior to presenting our sampling approach, we examine U-TAE on a standardized benchmark by adjusting sequence lengths and positional encodings to identify the minimal temporal resolution and ordering information needed for high classification accuracy. These insights directly inform our design choices and motivate the development of Thermal-Time-based Temporal Sampling (T3S), a model-agnostic method that aligns input selection with crop physiology.
@@
3.1 U-Net with Temporal Attention Encoder (U-TAE)
The U-TAE model is a spatio-temporal neural network designed to extract dense and temporally-aware features from satellite image time series. It consists of a shared 2D convolutional encoder applied to each timestamp independently, followed by a temporal self-attention module that aggregates the encoded features across time. This attention module uses positional encoding to preserve temporal ordering and applies scaled dot-product attention to dynamically weigh each timestamp based on its relevance. The aggregated multi-temporal representation is then decoded through a U-Net-style decoder with skip connections, enabling high-resolution semantic segmentation. The architecture can handle variable-length input sequences and produce temporally adaptive representations of land cover. Refer to Garnot and Landrieu, (2021) for more details.
@@ 3.2 Preliminary Analysis with U-TAE
  • Redundancy and noise in temporal observations. There is currently no comprehensive study establishing the exact number of time steps required to classify crops reliably using U-TAE or other architectures. Existing works vary widely, for example, Garnot and Landrieu, (2021) uses up to 61 observations, while Turkoglu et al., 2021a employs 71 timestamps. In principle, Sentinel-2 can provide an observation every 2 to 3 days over Switzerland, making very long sequences feasible. However, such long sequences are often redundant and noisy: Metzger et al., (2021) reports that roughly 50% of observations can be unusable due to cloud cover. Moreover, crops may develop relatively slowly, e.g. in periods with low temperatures. Consequently, analogous to a slowly evolving video, densely sampling in time adds little new information for distinguishing crop types.
  • Computational efficiency and memory constraints. Deep temporal models like U-TAE require substantial GPU memory when processing long sequences. To maintain a reasonable batch size during training, it is essential to limit the number of timestamps. Note that larger batches reduce the variance in the estimation of the gradient, resulting in a stronger and more stable training signal (Räisä et al.,, 2024). This is particularly advantageous for highly imbalanced datasets: choosing a batch size several times the minority-to-majority class ratio ensures that each minibatch contains sufficient minority-class examples to learn meaningful features (Google Developers,, 2025).
To ensure our evaluation remains unbiased, we base our initial analysis on the well-established PASTIS benchmark, a concise and reproducible toy crop classification dataset introduced alongside U-TAE by Garnot et al. (2021). By mirroring SwissCrop’s multi-temporal structure while remaining fully independent, PASTIS enables rapid iteration and guarantees that insights from our analysis directly inform algorithm design without the dataset itself biasing the method.
In this dataset each time series contains on average 48 Sentinel-2 observations (up to 61). We uniformly subsampled each sequence to 24 time steps, and trained with a batch size of 4 (see Garnot and Landrieu, (2021) for full training details). As shown in Table 1, this simple uniform subsampling yields only a 0.2% absolute drop in overall accuracy, despite no selective filtering of cloudy or low-information dates. Crucially, GPU memory usage is cut in half, and the number of multiply–accumulate operations (MACs) per forward pass with a single batch is reduced around 50%, dramatically lowering computational cost and freeing capacity for larger batches or faster training and inference. Accordingly, we set the temporal length to 24 for the experiments.
Table 1: Performance comparison for varying temporal length on the Pastis dataset. Results are averaged over a 5-fold experimental setting of the Pastis dataset. Refer to Garnot and Landrieu, (2021) for details. M stands for GPU memory.
@@
  • calendarPE: Positions are encoded as the actual calendar dates of each observation (1, …, 365).
@@
In summary, our preliminary findings reveal that (i) substantial temporal subsampling incurs only a marginal accuracy loss while reducing both GPU memory usage and computational cost drastically, and (ii) U-TAE relies chiefly on the relative ordering of observations, whereas absolute timestamps offer limited benefit. These insights suggest that, rather than encoding raw dates, we should focus on selecting the most informative, developmentally relevant observations. In the next section, we leverage this principle to introduce model‐agnostic approach that partitions input sequences according to growing degree days to better align sampling with crop physiology.

### v3

@@ Abstract
Crop type classification using optical satellite time series remains limited in its ability to generalize across seasons, particularly when crop phenology shifts due to inter-annual weather variability. This hampers real-world applicability in scenarios where current-year labels are unavailable. In addition, uncertainty quantification is often overlooked, which reduces the reliability of such approaches for operational crop monitoring. Inspired by ecophysiological principles of plant growth, we propose a simple, model-agnostic Thermal-Time-based Temporal Sampling (T3S) method that replaces calendar time with thermal time. By subsampling time series in this biologically meaningful way, our method highlights key periods within the growing season while reducing temporal redundancy and noise. We evaluate the T3S on a multi-year Sentinel-2 dataset covering the entirety of Switzerland, which allows us to assess all applied methods on unseen years. Compared to state-of-the-art baselines, our approach yields substantial improvements in classification accuracy and, critically, provides well-calibrated uncertainty estimates. Moreover, the T3S method excels in low-data regimes and enables significantly more accurate early-season classification. With just 10% of the training labels, it outperforms the current baseline in both accuracy and uncertainty calibration, and by the end of June, it achieves a performance similar to the full-season baseline model. The dataset and the code will be publicly available.
@@ 1 Introduction
Monitoring agricultural land use is crucial for the sustainability of our food systems [16]. As the world’s population grows, climates becomes extremer, and diets evolve, there is a pressing need for monitoring tools that deliver reliable, affordable, and scalable insights into farming practices. In contrast to traditional methods, such as field surveys and self-reported data, which can be labor-intensive, expensive, and susceptible to errors, remote sensing-based approaches offer a faster, more cost-effective, and less subjective alternative for tracking land-use dynamics [58].
Recent progress in remote sensing technologies and machine learning algorithms has significantly enhanced the ability to monitor agricultural land use. High-resolution, multi-spectral data from platforms such as Sentinel-2 provide frequent, detailed observations of croplands at minimal cost. By applying advanced machine learning methods, including conventional ensemble learners, e.g, XGBoost [8] and deep neural networks, e.g, Transformers [56], to these time series images, we can capture crop-specific growth dynamics and distinguish among diverse management practices with high precision. These integrated approaches enable automated, scalable mapping of crop types and growth stages, supporting timely, data driven decisions for sustainable agriculture [50].
Figure 1: Spatially averaged cumulative growing degree days (cGDD) across Switzerland for multiple years.
Modern deep learning methods [14, 51, 29, 48, 40] excel in classifying crop types from optical satellite image time series. However, these models are typically trained on curated datasets with limited size and/or validated on data collected within a single growing season [39, 42, 46, 14, 51, 35], neglecting interannual shifts in growth driven by climate variability and thus limiting their applicability across years and regions. Furthermore, their dependence on current-year labeled data for training precludes real time deployment during the ongoing season. As also noted by [37], model development should focus on datasets with real-world relevance to further improve accuracy, generalization ability, and real impact on end user applications. Lastly, although these approaches may achieve state-of-the-art accuracy, they rarely incorporate uncertainty quantification, leaving predictive confidence unassessed and often uncalibrated. Ensuring the models are well-calibrated makes them reliable components of agricultural monitoring systems, where miscalibration could otherwise lead to costly resource misallocations or risk mismanagement.
Figure 2: NDVI time series of two parcels with the same crop rotation over two years (2021 and 2022), aligned by calendar date (left) and by cumulative growing degree days (cGDD) (right).
@@
In this work, we build upon the state-of-the-art attention-based deep learning architecture proposed by Garnot and Landrieu [14], and introduce a simple, model-agnostic sampling strategy that incorporates daily average temperature to reduce redundancy and noise in satellite image time series. Deep learning models for crop classification typically rely on subsampled time series, as crop development progresses gradually and many satellite observations are temporally redundant. Subsampling is also necessary to reduce sequence length for efficient model training and inference within limited GPU memory. To better guide this sampling, we reparameterize calendar time into thermal time, using cumulative growing degree days (GDD) as our temporal reference, and then perform uniform subsampling in this thermal‐time domain. This ensures the model concentrates on the crop’s biologically active periods, focusing on the most informative observations.
@@ 2 Related Work
Deep learning has revolutionized crop‐type classification from satellite time series. Early sequence models based on LSTM [38] and convolutional‐RNN architectures [51] harnessed phenological dynamics, while attention‐based approaches such as self‐attention [41, 15], the lightweight Temporal Attention Encoder (L‐TAE) [44], and U‐Net–integrated attention [14] improved long‐range dependencies. Vision Transformer adaptations [2] and state‐space models like Mamba [33] further enhanced spatio‐temporal representations. Integrating domain knowledge, convSTAR [55] and prototype‐based supervision [45] has been used to incorporate class relationships. The most relevant method for crop type mapping is thermal-time positional encoding [31], which aligns phenological stages across regions using cumulative growing degree days. To handle irregular sampling and early‐season needs, methods incorporating reward‐based early classification [43], neural ODEs [29], and adaptive sampling frameworks [7] have been developed. Generalizing across seasons is hampered by interannual phenological shifts and sensor inconsistencies. Multi‐year evaluations [54, 5] and methods such as Fourier‐based time series reconstruction with gradient boosting [9], Google Earth Engine workflows [32], and photometric augmentation in U‐Net models [22] have shown promising improvements. Uncertainty estimation enhances trust in crop maps. Deep ensembles [23] remain the standard, with efficient approximations including MC Dropout [12], Snapshot Ensembles [20], Masksembles [10], FiLM‐Ensemble [53], and LoRA‐Ensemble [19]. In satellite‐based crop monitoring, stochastic inference quantifies per‐pixel and per‐field uncertainty, guiding data collection and decision‐making. MC Dropout has been applied to winter wheat yield forecasting [49], annual cropland mask segmentation [22], and ViT–ResNet feature encoding for field‐level confidence scoring [27]. Probabilistic deep ensembles trained on multi‐sensor inputs produce robust uncertainty distributions, as demonstrated in German crop‐type mapping [3]. Integrating these uncertainty estimates with deterministic predictions flags low‐confidence areas, enabling targeted ground truthing and more reliable operational deployment of crop‐type mapping systems. Refer to Appendix C for an extended version of the Related Work section.
@@ 3 Method
Formally, the objective is to predict a crop type map $\bm{y}\in{\rm I\!R}^{H\times W\times C}$ from a sequence of input images $\bm{X}=\{\bm{x}_{1},\bm{x}_{2},..,\bm{x}_{L}\}\in{\rm I\!R}^{H\times W\times B}$. $H$ and $W$ are the height and width of the input images, respectively, $B$ is the number of input bands, $L$ is the number of time stamps in the input sequence that can differ from sample to sample, and $C$ is the number of crop types. Our method builds on the U-Net with Temporal Attention Encoder (U-TAE) [36, 14], a state-of-the-art deep learning model for dense segmentation of satellite image time series. Prior to presenting our sampling approach, we examine U-TAE on a standardized benchmark by adjusting sequence lengths and positional encodings to identify the minimal temporal resolution and ordering information needed for high classification accuracy. These insights directly inform our design choices and motivate the development of Thermal-Time-based Temporal Sampling (T3S), a model-agnostic method that aligns input selection with crop physiology.
@@
3.1 U-Net w/ Temporal Attention Encoder (U-TAE)
The U-TAE model fuses spatial feature extraction with temporal attention to segment land cover in satellite image time series $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L}\}$. Each frame $\mathbf{x}_{i}$ is passed through a shared 2D convolutional encoder, producing per-frame feature maps. These maps are then aggregated by a self-attention block that employs positional encodings and scaled dot-product attention to weight timestamps according to their phenological relevance. The resulting multi-temporal embedding is fed into a U-Net–style decoder with skip connections, yielding high-resolution, temporally adaptive segmentation masks. U-TAE naturally handles sequences of varying length and is optimized via pixel-wise cross-entropy loss against the ground-truth labels. For training details, refer to the Appendix A.
@@ 3.2 Preliminary Analysis with U-TAE
  • Redundancy and noise in temporal observations. There is currently no comprehensive study establishing the exact number of time steps required to classify crops reliably using U-TAE or other architectures. Existing works vary widely, for example, Garnot and Landrieu [14] uses up to 61 observations, while Turkoglu et al. [51] employs 71 timestamps. In principle, Sentinel-2 can provide an observation every 2 to 3 days over Switzerland, making very long sequences feasible. However, such long sequences are often redundant and noisy: Metzger et al. [29] reports that roughly 50% of observations can be unusable due to cloud cover. Moreover, crops may develop relatively slowly, e.g. in periods with low temperatures. Consequently, analogous to a slowly evolving video, densely sampling in time adds little new information for distinguishing crop types.
  • Computational efficiency and memory constraints. Deep temporal models like U-TAE require substantial GPU memory when processing long sequences. To maintain a reasonable batch size during training, it is essential to limit the number of timestamps. Note that larger batches reduce the variance in the estimation of the gradient, resulting in a stronger and more stable training signal [34]. This is particularly advantageous for highly imbalanced datasets: choosing a batch size several times the minority-to-majority class ratio ensures that each minibatch contains sufficient minority-class examples to learn meaningful features [17].
To ensure our evaluation remains unbiased, we base our initial analysis on the commonly used, well-established PASTIS benchmark, a concise and reproducible toy crop classification dataset introduced alongside U-TAE by [14]. By mirroring SwissCrop’s multi-temporal structure while remaining fully independent, PASTIS enables rapid iteration and guarantees that insights from our analysis directly inform algorithm design without the dataset itself biasing the method.
In this dataset each time series contains on average 48 Sentinel-2 observations (up to 61). We uniformly subsampled each sequence to 24 time steps, and trained with a batch size of 4 (see [14] for full training details). As shown in Table 1, this simple uniform subsampling yields only a 0.2% absolute drop in overall accuracy, despite no selective filtering of cloudy or low-information dates. Crucially, GPU memory usage is cut in half, and the number of multiply–accumulate operations (MACs) per forward pass with a single batch is reduced around 50%, dramatically lowering computational cost and freeing capacity for larger batches or faster training and inference. Accordingly, we set the temporal length $T=24$ for the experiments.
Table 1: Performance comparison for varying temporal length on the Pastis dataset. Results are averaged over a 5-fold experimental setting. Refer to [14] for details. M stands for GPU memory.
@@
  • calendarPE: Positions are encoded as the actual calendar dates of each observation, e.g., (2, 5,…, 365).
@@
In summary, our preliminary findings reveal that (i) substantial temporal subsampling incurs only a marginal accuracy loss while drastically reducing both GPU memory usage and computational cost, and (ii) U-TAE relies chiefly on the relative ordering of observations, whereas absolute timestamps offer limited benefit. These insights suggest that, rather than encoding raw dates, we should focus on selecting the most informative, developmentally relevant observations. In the next section, we leverage this principle to introduce model‐agnostic approach that partitions input sequences according to growing degree days to better align sampling with crop physiology.